# AgentDR: Dynamic Recommendation with Implicit Item-Item Relations via LLM-based Agents

### v1

@@ 4.3. RQ1: Ranking Performance Comparison
Dataset Instacart Electronics Sports
Metric R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20
@@ 4.4. RQ2: VDCG Comparison
Table 4. Performance comparison on the proposed VDCG which jointly evaluate the semantic and ordering correctness.
Dataset Instacart Electronics Sports
Metric V@5 V@10 V@5 V@10 V@5 V@10
@@ 4.5.2. Results
Dataset Instacart Electronics Sports
Metric N@10 N@20 N@10 N@20 N@10 N@20
@@ 4.6. RQ 4: Ablation Study
To quantify the contribution of each LLM-based module, we compare the performance of AgentDR with different modules removed or replaced. For fair comparison, we use the original ranking comparison as quantitative comparison module.
@@ 5. Conclusion
In this work, we propose a novel LLM-based agent framework, AgentDR, for full-ranking recommendation. The commonsense world knowledge of the LLM is utilized to infer user intent regarding substitutes and complements. We acknowledge that a limitation of this work is that substitution and complementarity relationships are more prevalent in domains like groceries than in others such as movies or music. Future work may explore knowledge sharing among user agents to better capture global data characteristics tailored to specific downstream scenarios.
@@ A.1. Implementation Details
Since most item titles in the three grocery datasets used in experiments are self-explainable, we use titles as textual descriptions of items and remove all the items without titles. For each dataset, users’ interactions are chronologically organized based on timestamps to construct user behavior sequences. Since our framework delegates full-ranking sequential recommendation tasks to existing models, we follow the same data preprocessing strategy in previous works to remove cold-start users and items (He et al., 2020; Kang and McAuley, 2018; Mao et al., 2021). The recommendation tools are pretrained on the first (nk)(n-k) items of each user’s interaction sequence $\mathbf{s}_{[:-k]}$, following the standard training protocol in recommendation where models learn from early interactions and are evaluated on their ability to predict future behaviors. This setup ensures consistency across tools and enables fair performance comparison between our framework and other recommendation baselines. A grid search is applied for searching learning rates $\alpha,\beta$ and $\gamma$ in {1e-3, 5e-3, 1e-2, 5e-2, 1e-1}. To mitigate overfitting, each learning rate is multiplied by a decay factor of 0.8 after every optimization epoch. The number of historical items $c$ used for substitute and complement generation is set to 10. Following the training scheme in sequential recommendation, the number of target items $k$ is set to 1 during optimization. Since the evaluation metrics consider at most top 20 items, the number of reranking candidates $k^{\prime}$ is fixed at 20. All experiments are conducted using the Phi-4 language model, deployed locally with vLLM on eight Tesla V100 GPUs, each with 32 GB of VRAM. The temperature is set to 0 to ensure deterministic outputs. Aside from the complexity of the recommendation tools described in Sec. 3.1.2, the quantized version of a single Phi-4 model requires less than 12 GB of VRAM. Following prior LLM-agent works (Zhang et al., 2024b; Wang et al., 2025b; Huang et al., 2025b), we train the recommendation tools on the entire training set, and randomly sample 160 agents for optimization on each dataset.
A.2. Agent Memory Visualization
@@
A.3. Details of Prompts
The detailed contents of prompts used in AgentDR are shownd from Table 7 to Table 13.
@@ A.3. Details of Prompts
A.4. Additional Related Works
A.4.1. LLM for Recommendation
@@
The capabilities of traditional embedding-based recommender systems remain limited when relying solely on historical interaction data. To address this limitation, prior works have incorporated LLMs either as text encoders to enrich recommendation embeddings with textual information (Wei et al., 2024; Ren et al., 2024) or as standalone recommenders by reformulating recommendation tasks as language modeling problems (Geng et al., 2022; Cui et al., 2022). While these approaches leverage the world knowledge of pretrained language models and show advantages in cold-start scenarios (Huang et al., 2025a; Hou et al., 2024), their effectiveness is constrained by the fixed dimensionality of feature vectors and the fundamental mismatch between language modeling and behavior modeling, which hinders their ability to fully exploit interaction signals. We compare these pioneer works with LLM-based recommendation agents (Zhang et al., 2024b, a; Xu et al., 2025; Huang et al., 2025b; Wang et al., 2025b, 2024) in Table 1.
@@
To tightly couple the pretrained knowledge in LLMs and ID-based behavior patterns in recommender systems, recent works introduce user and item tokens into the LLM vocab space, and align these tokens with the language space (Liao et al., 2024; Zhu et al., 2024). However, compared to inference-only LLM agents for recommendation, the effectiveness of this knowledge alignment with LLMs is sensitive to the amount of tunable parameters and available data. Besides, the concept of Retrieval-Augmented Generation (RAG) has been adapted for recommendation, where candidate interactions are first retrieved without LLMs and then reasoned by LLMs (Wang et al., 2025a; Wu et al., 2024a). Retrieval alleviates the reasoning load for LLMs by narrowing the candidate space, but the overall performance is influenced by the inherent biases of the specific deployed retrieval method. In contrast, LLM agents usually leverage multiple tools to enhance the performance on the target task, like AgentDR equipped with full-ranking tools for recommendation.
A.4.2. Substitution and Complementarity
@@
Substitution and complementarity as item-item relations play a crucial role in understanding user intent. Substitutes help identify alternatives when a preferred item is unavailable or when the user seeks variety, while complements reveal co-purchase patterns that often reflect planned consumption behaviors. However, although ground-truth labels for substitutes or complements benefit recommendation (Zhang et al., 2022; Wu et al., 2023a), they are typically unavailable in most public datasets (Ni et al., 2019; Bonab et al., 2021a). As a result, previous works (Chen et al., 2020a; Liu et al., 2020; Xu et al., 2020) constructed these labels from co-view and co-purchase logs of users due to the inefficiency of rigorous label annotation by human annotator. However, many item-pairs labeled based on these behavior logs are not functionally substitution or complementary (Jian et al., 2022; Ye et al., 2023; Sugahara et al., 2024). To address this gap, LLMs have been applied for labeling substitution and complementary relations in previous works (Papso, 2023; Hosseini et al., 2025), indicating that their extensive world knowledge make LLMs particularly well-suited for reasoning over these implicit item-item relationships in e-commerce datasets. In AgentDR, these relationships are leveraged during intent discrimination and relational reranking to enhance recommendation relevance.

### v2

@@ 4.3. RQ1: Ranking Performance Comparison
Instacart Electronics Sports
R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20
@@ 4.4. RQ2: VDCG Comparison
Table 4. Performance comparison on VDCG, which jointly evaluates semantic and ordering correctness.
Instacart Electronics Sports
V@5 V@10 V@5 V@10 V@5 V@10
@@ 4.5.2. Results
Instacart Electronics Sports
N@10 N@20 N@10 N@20 N@10 N@20
@@ 5. Conclusion
In this work, we propose a novel LLM-based agent framework, AgentDR, for full-ranking recommendation. The commonsense world knowledge of the LLM is utilized to infer user intent regarding substitutes and complements. A limitation of this work is that substitution and complementarity relationships are more prevalent in domains like groceries than in others such as movies or music.
Acknowledgements
This work is supported in part by NSF under grants III-2106758, and POSE-2346158.
@@ A.1. Implementation Details
Since most item titles in the three grocery datasets used in experiments are self-explainable, we use titles as textual descriptions of items and remove all the items without titles. For each dataset, users’ interactions are chronologically organized based on timestamps to construct user behavior sequences. Since our framework delegates full-ranking sequential recommendation tasks to existing models, we follow the same data preprocessing strategy in previous works (He et al., 2020; Kang and McAuley, 2018; Mao et al., 2021) to remove cold-start users and items. The recommendation tools are pretrained on the first (nk)(n-k) items of each user’s interaction sequence $\mathbf{s}_{[:-k]}$, following the standard training protocol in recommendation where models learn from early interactions and are evaluated on their ability to predict future behaviors. This setup ensures consistency across tools and enables fair performance comparison between our framework and other recommendation baselines. A grid search is applied for searching learning rates $\alpha,\beta$ and $\gamma$ in {1e-3, 5e-3, 1e-2, 5e-2, 1e-1}. To mitigate overfitting, each learning rate is multiplied by a decay factor of 0.8 after every optimization epoch. The number of historical items $c$ used for substitute and complement generation is set to 10. Following the training scheme in sequential recommendation, the number of target items $k$ is set to 1 during optimization. Since the evaluation metrics consider at most top 20 items, the number of reranking candidates $k^{\prime}$ is fixed at 20. All experiments are conducted using the Phi-4 language model, deployed locally with vLLM on eight Tesla V100 GPUs, each with 32 GB of VRAM. The temperature is set to 0 to ensure deterministic outputs. Aside from the complexity of the recommendation tools described in Sec. 3.1.2, the quantized version of a single Phi-4 model requires less than 12 GB of VRAM. Following prior LLM-agent works (Zhang et al., 2024b; Wang et al., 2025a; Huang et al., 2025c), we train the recommendation tools on the entire training set, and randomly sample 160 agents for optimization on each dataset.
A.2. Cost Analysis
The total LLM API calls in AgentDR consist of three parts: generation, agent optimization, and reranking. Generation (Eq. 1, Eq. 2 and Eq. 3) and reranking (Eq. 10, Eq. 12, and Eq. 13) contain 6 calls per user. Each epoch in agent optimization contains 3 calls (Eq. 4, Eq. 8 and Eq. 11) per user. Hence, the total number of calls is $6n+3nt$ where $n$ and $t$ are the numbers of users and epochs, respectively. With all the agent memory updated, AgentDR takes less than 1 second to generate the final ranking list for each user. We acknowledge that LLM-based agents generally have higher inference latency than embedding-based recommendation models, but they are still practical for offline ranking or secondary-stage recommendation pipelines.
A.3. Agent Memory Visualization
@@
A.4. Details of Prompts
The detailed contents of prompts used in AgentDR are shown from Table 7 to Table 13. The prompts for VDCG evaluation is shown in Table 14.
@@ A.4. Details of Prompts
[Instruction]
Given the same category, rate how well each item from the recommended list matches the target item based on relevance, usefulness, and user interest.
[Category]
{item_category}
[Target Item]
{item_description}
[Recommended List]
{item_descriptions}
The output must be a list of integers. Each score must be between 0 (very unrelated) and 9 (exact match).
Table 14. Prompts for the LLM to rate how well each item in the recommended list aligns with the ground-truth item for VDCG evaluation in Sec. 4.2.
A.5. Additional Related Work
In this appendix, we discuss additional works related to ours, including research on LLM-based recommendation systems and methods for identifying substitute and complementary products.
A.5.1. LLM for Recommendation
@@
The capabilities of traditional embedding-based recommender systems remain limited when relying solely on historical interaction data. To address this limitation, prior works have incorporated LLMs either as text encoders to enrich recommendation embeddings with textual information (Wei et al., 2024; Ren et al., 2024; Yang et al., 2025b, 2024a; Zhu et al., 2025; Nerrise et al., 2025) or as standalone recommenders by reformulating recommendation tasks as language modeling problems (Geng et al., 2022; Cui et al., 2022; He et al., 2025). While these approaches leverage the world knowledge of LLMs and show advantages in cold-start scenarios (Huang et al., 2025a; Hou et al., 2024), their effectiveness is constrained by the fixed dimensionality of feature vectors and the fundamental mismatch between language and behavior modeling, which hinders their ability to fully exploit interaction signals. We compare these pioneer works with LLM-based recommendation agents (Zhang et al., 2024b, a; Xu et al., 2025; Huang et al., 2025c; Wang et al., 2025a, 2024) in Table 1.
@@
To tightly couple the pretrained knowledge in LLMs and ID-based behavior patterns in recommender systems, recent works introduce user and item tokens into the LLM vocab space, and align these tokens with the language space (Liao et al., 2024; Zhu et al., 2024; Yang et al., 2025a). However, compared to inference-only LLM agents for recommendation, the effectiveness of this knowledge alignment with LLMs is sensitive to the amount of tunable parameters and available data. Besides, the concept of Retrieval-Augmented Generation (RAG) has been adapted for recommendation, where candidate interactions are first retrieved without LLMs and then reasoned by LLMs (Wang et al., 2025b; Wu et al., 2024a). Retrieval alleviates the reasoning load for LLMs by narrowing the candidate space, but the overall performance is influenced by the inherent biases of the specific deployed retrieval method. In contrast, LLM agents usually leverage multiple tools to enhance the performance on the target task, like AgentDR equipped with full-ranking tools for recommendation.
A.5.2. Substitution and Complementarity
@@
Although substitution and complementarity as item-item relations play a crucial role in understanding user intent (Zhang et al., 2022; Wu et al., 2023; Nerrise et al., 2025), they are unavailable in most public datasets (Ni et al., 2019; Bonab et al., 2021b). As a result, previous works (Chen et al., 2020b; Liu et al., 2020; Xu et al., 2020) constructed these labels from co-view and co-purchase logs of users due to the inefficiency of rigorous label annotation by human annotator. However, many item-pairs labeled based on these behavior logs are not functionally substitution or complementary (Jian et al., 2022; Ye et al., 2023; Sugahara et al., 2024). To address this gap, LLMs have been applied for labeling substitution and complementary relations in previous works (Papso, 2023; Hosseini et al., 2025), indicating that their extensive world knowledge make LLMs particularly well-suited for reasoning over these implicit item-item relationships in e-commerce datasets. In AgentDR, these relationships are leveraged during intent discrimination and relational reranking to enhance recommendation relevance.